1) A 141 g sample of carbon from ancient wood is found to have an activity of 5.289 decays/min. a. How many half lives have elapsed since the tree was cut down? b. How many years have elapsed since the tree was cut down?

would appreciate any help
Answer #1

apply activity = dN/dt = LN

L = 5.289/60/141

L = 6.251*10^-4

as decay consatnt L = Ln2/T1/2

T/12 = 0.693/6.251 e-4

T/12 = 1108.48 years--------------answer to part B

------------------------------------------

apply N = N0 e^-Lt

e^-Lt = 1/141) = 0.00709

-Lt = 4.94

t = 4.94/6.251 e-4 = 7916 .67 Years------answers A
Answer #2

decay constant (?): 5.289

half-life (t1/2): 0.13105448678389
mean lifetime (?): 0.18907165815844